1.Mark decides to go swimming in the river to cool off Mark is able to swim at 2.6 m/s in calm water. He aims straight South (River runs east west) across the river which has a current of 1.1 m/s East. 

a. Determine Mark’s velocity as seen from the river bank.

b. If it takes him 23 seconds to travel across the river, how far downstream did he end up after his swim


1
Expert's answer
2021-06-21T11:38:47-0400

a)

"v'=\sqrt{v^2+u^2},"

"v'=\sqrt{2.6^2+1.1^2}=2.8~\frac ms,"

b)

"l=ut,"

"l=1.1\cdot 23=2.5~m."